# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 5.1.1. Execution Environment
Virtual machines (VMs) offer a layer of abstraction over physical hardware, allowing multiple independent operating systems to share resources efficiently (Sharma et al., 2016; Smith and Nair, 2005). However, VMs can experience performance issues when resources are overutilized, and they are vulnerable to hypervisor failures and migration difficulties (Azure Readiness, 2023; Fogli et al., 2023). Chaos engineering addresses these challenges by testing VMs under stress conditions such as CPU or memory exhaustion to assess if VMs can recover from failures or seamlessly migrate workloads (Kostic, 2024; Chen et al., 2024; Green, 2023).
Containers, which encapsulate applications and their dependencies in lightweight environments, provide greater efficiency and portability compared to VMs (Sharma et al., 2016; Koskinen et al., 2019). However, containers face resource contention, network instability, and orchestration failures, especially in platforms such as Docker and Kubernetes (Watada et al., 2019). Chaos engineering tests on containers simulate abrupt stoppages and network disruptions, helping to evaluate containers’ ability to recover within container orchestration systems while maintaining application performance  (Simonsson et al., 2021; Camacho et al., 2022; Bremmers, 2021; Kostic, 2024; Ikeuchi et al., 2020).
Serverless functions allow applications to be triggered by specific events and run in isolated containers without requiring infrastructure management (Castro et al., 2017; Taibi et al., 2020). While serverless architectures are scalable and cost-effective, they face challenges such as resource limitations and the reliance on third-party cloud providers (Rajan, 2020; van Eyk and Iosup, 2018; Kadikar, 2023; Camacho et al., 2022). Chaos engineering helps ensure that serverless functions can handle these challenges by simulating cold starts and execution failures, verifying that functions recover quickly and maintain performance under varying loads (Starr et al., 2022; Jernberg et al., 2020; Gill, 2022; Bedoya et al., 2023).
@@
Public cloud platforms, such as AWS, Microsoft Azure, and Google Cloud, offer computing resources over the Internet that are shared between multiple organizations (Saraswat and Tripathi, 2020). This model enables businesses to scale their infrastructure on demand, eliminating the need for significant investments in physical hardware. Due to its cost efficiency and scalability, the public cloud has become a popular choice for many enterprises (Dutta and Dutta, 2019). However, relying on shared infrastructure introduces specific vulnerabilities. For example, a regional outage or failure in an availability zone can have widespread impacts, simultaneously affecting many customers (Paavolainen, 2016). In addition, many organizations depend on third-party services, such as application programming interfaces (APIs) or managed databases, and failures in these services can trigger further disruptions in their systems (Zhou et al., 2010). To address these risks, chaos engineering can simulate large-scale failures, such as regional outages or availability zone failures, to test how well systems can failover to other regions without significant service interruptions (Fogli et al., 2023). Moreover, by injecting failures into critical third-party services, chaos engineering enables organizations to evaluate the effectiveness of their recovery strategies, ensuring that operations can continue even in the face of service disruptions (Ji et al., 2023; Yu et al., 2021).
@@
Private cloud environments, which provide dedicated infrastructure exclusively for a single organization, can be hosted on-premises or managed by third-party providers, offering greater control over data privacy, security, and customization (Jadeja and Modi, 2012; Gastermann et al., 2015). This model suits organizations with strict regulatory requirements or managing sensitive information (Pearson, 2013). Despite their benefits, private clouds often face scalability and geographic redundancy challenges, as they typically lack the built-in redundancy of public clouds, making them more susceptible to service interruptions due to hardware failures or resource constraints (Kumar and Goudar, 2012). Without adequate failover mechanisms, private cloud environments may struggle to maintain operational continuity during hardware or network failures (Rimal et al., 2009; Tivey et al., 2019). Chaos engineering can be used in private clouds to simulate hardware failures, such as server crashes or network partitions, to evaluate the system’s ability to recover from localized disruptions (Patel, 2022). By simulating resource exhaustion, such as CPU or memory overloads, organizations can test whether their infrastructure can scale to meet unexpected surges in demand, ensuring continuous service availability under pressure (Gill, 2021).
@@
Finally, hybrid cloud environments combine the strengths of public and private clouds, allowing organizations to manage sensitive workloads on private infrastructure while leveraging the scalability of public clouds for nonsensitive or high-demand tasks (Goyal, 2014). This approach provides flexibility but introduces complexity in maintaining smooth integration between the two environments (Lu et al., 2014). A key challenge of hybrid clouds is to ensure seamless communication between public and private components, as issues such as network latency, connectivity problems, and data synchronization delays can arise when workloads are distributed across both environments, potentially leading to service disruptions (Annapureddy, 2010). Chaos engineering can address this issue by allowing organizations to simulate network interruptions or increased latency between public and private clouds (Fogli et al., 2023). These tests help to evaluate whether workloads can continue to operate despite connectivity issues (Torkura et al., 2020). Furthermore, chaos experiments that simulate delays or failures in data replication help ensure that the system maintains data consistency and operational continuity, even when part of the infrastructure is compromised (Camacho et al., 2022).
@@
This refers to the computing infrastructure hosted and managed within an organization’s physical location rather than relying on cloud providers (Willard and Hutson, 2024; Nikolaidis et al., 2021; Väisänen, 2023). These systems offer greater control over hardware, data, and network resources, essential for industries with strict regulatory requirements. However, they also face hardware failures and scaling limitations, affecting the reliability of the systems (Byström, 2022). For example, an organization managing critical on-premise data might simulate a hardware failure, such as a server or disk crash, to test how quickly backup systems respond. If synchronization delays are found, the replication process can be optimized to ensure smooth transitions and compliance with data availability standards (Fogli et al., 2023; Dedousis et al., 2023). Similarly, scaling limitations can be addressed by testing how an on-premise infrastructure responds to various demand surges, such as a sudden spike in user traffic during an e-commerce company’s high-demand event (Azure Readiness, 2023; Gianchandani et al., 2022). This helps identify bottlenecks, optimize resource allocation, and ensure that the system remains responsive under varying load conditions (Naqvi et al., 2022; Wu, 2021; Sbiai et al., 2023; Wickramasinghe, 2023).
@@ 5.1.2. Experimentation Modes
Manual experimentation is most appropriate when engineers require high precision and control, as they take full responsibility for designing, configuring, and executing chaos tests (Kostic, 2024). This approach is suited for organizations with little or no experience in chaos engineering or when highly specific, customized experiments are necessary (Durai, 2022). For example, an engineer might manually trigger a database failure or introduce network latency to observe the system response in real time (Jernberg et al., 2020; Bremmers, 2021). While this mode allows for fine-tuned control, it can be time-consuming and prone to human error, making it ideal for small-scale, customized tests or critical systems where close monitoring is essential (Ikeuchi et al., 2020; Chen et al., 2024).
Semi-automated experimentation provides a middle ground between manual control and complete automation, where engineers define the experiment parameters (Sondhi et al., 2021; Jernberg et al., 2020). At the same time, the platform automates tasks such as triggering tests and collecting data, making this mode ideal for experiments that benefit from automated execution while still requiring moderate human involvement to reduce manual effort (Kesim et al., 2020; Nikolaidis et al., 2021; Gill, 2021). For example, engineers might set up an experiment to test resource exhaustion on specific microservices, with the platform automatically running the test at predefined intervals, collecting performance metrics, and allowing them to focus on adjusting parameters and interpreting results (Pierce et al., 2021; Dedousis et al., 2023).
Fully automated experimentation suits large-scale systems that require frequent and continuous testing (Zhang et al., 2021a; Simonsson et al., 2021; Basiri et al., 2019). After engineers configure the initial setup, the chaos engineering platform takes over, seamlessly integrating experiments into the system operations, typically through the CI/CD pipelines (Konstantinou et al., 2021; Pulcinelli et al., 2023; Manna, 2021; Colyer, 2019; Azure Readiness, 2023). For example, after a code deployment or infrastructure update, the platform can automatically simulate a service outage or introduce network disruptions, then autonomously gather performance metrics, identify system weaknesses, and generate detailed reports, requiring engineers to intervene only when necessary (Dedousis et al., 2023; Bairyev, 2023; Kesim et al., 2020; Cotroneo et al., 2022; Nikolaidis et al., 2021; Patel, 2022).
@@ 5.1.3. Chaos Automation Strategy
This involves developing programs or scripts to extend the capabilities of existing tools and gain precise control over system configurations and experiments (Myers et al., 2000; Thüm et al., 2014; Khankhoje, 2022). It allows engineers to simulate unique failure conditions, ranging from simple component failures, such as restarting a single service, to complex multi-step sequences involving multiple systems, tailored to the specific challenges of their architecture that standard tools may not address (Gill, 2021; Inc., 2023; Torkura et al., 2020; Azure Readiness, 2023; Patel, 2022; Nikolaidis et al., 2023). For example, custom code can terminate random containers in distributed environments such as Kubernetes, simulating unexpected service failures and testing the system’s recovery capacity by automatically restarting services (Bremmers, 2021; Fawcett, 2020; Simonsson et al., 2021; Vu et al., 2022). In more complex scenarios, custom code can initiate a service outage followed by a database failure or a network partition to observe how these failures impact system recovery and resilience (Wickramasinghe, 2023; Gill, 2021; Simonsson et al., 2021). Additionally, custom code can be integrated into CI/CD pipelines to simulate high resource consumption, such as CPU throttling or memory exhaustion, after each deployment, allowing early detection of performance bottlenecks before they affect production (Azure Readiness, 2023; Konstantinou et al., 2021; Manna, 2021; Kadikar, 2023; Palani and Gupta, 2023). This custom code for chaos experiments can be written in various programming languages, such as Python, Bash, or Go, depending on the complexity and specific requirements of the experiment (Siwach et al., 2022; Malik et al., 2023; Sondhi et al., 2021; Simonsson et al., 2021; Zhang et al., 2021b).
@@
This process automatically integrates external systems, such as monitoring, log, observability, alerting, incident management, and security platforms, into chaos engineering experiments (Torkura et al., 2020; Simonsson et al., 2021). This automated integration ensures continuous data collection without manual intervention, enabling real-time analysis during system disruptions or failures. As a result, teams can monitor performance, identify issues quickly, and respond more efficiently to failures (Palani and Gupta, 2023; Kostic, 2024). This process is vital to integrating monitoring tools that track essential system metrics such as CPU usage, memory consumption, and network performance (Abdul, 2024; Wu, 2021; Gianchandani et al., 2022; Zhang et al., 2019). Monitoring tools can automatically detect performance deviations and provide immediate feedback. For example, if a service is overloaded due to an unexpected spike in user requests, monitoring systems measure the increase in CPU usage and network load, helping the team assess the impact and determine whether the system can handle the demand or if further intervention is needed (Wickramasinghe, 2023; Fogli et al., 2023; Cotroneo et al., 2022; Gill, 2021; Siwach et al., 2022; Malik et al., 2023). Logging tools record vital events such as service failures, error messages, and recovery attempts, documenting the system’s response in real-time. For example, if network communication between two microservices fails, the logging systems capture the exact moment the failure occurred and show the error messages generated by failed retries (Cotroneo et al., 2022; Ma et al., 2023). This data helps teams analyze how the failure affected communication and whether recovery mechanisms worked effectively (Naqvi et al., 2022; Malik et al., 2023; Kostic, 2024; Palani and Gupta, 2023). Building on this, observability tools provide a broader view by correlating logs and metrics to show how different components of the system interact during failure scenarios. These tools enable teams to identify how failures propagate across interconnected services (Simonsson et al., 2021; Manna, 2021; Green, 2023; Kostic, 2024; Krishnamurthy, 2021). For example, if latency is introduced between services in a chaos experiment, observability systems track the delays at each hop between services and how this latency affects downstream components such as database queries or API responses. This tracing allows the team to see if minor delays in one service cause issues elsewhere (Durai, 2022; Katirtzis, 2022; Kostic, 2024; Chen et al., 2024; Nikolaidis et al., 2023; Frank et al., 2023). Furthermore, alert notification systems can ensure that relevant teams are promptly notified when critical failures occur or performance thresholds are exceeded (Tivey et al., 2019; Gianchandani et al., 2022; Palani and Gupta, 2023; Bremmers, 2021; Starr et al., 2022). For example, if a database connection pool reaches its maximum capacity during a high-load chaos test, the alert system triggers a notification to the database operations team, allowing them to address the issue before it causes a system-wide failure. This early warning helps prevent more significant disturbances (Green, 2023; Malik et al., 2023). Finally, incident management tools streamline the resolution process in conjunction with alerting by automatically generating tickets for failures detected during chaos tests. These tickets are assigned to the appropriate teams and tracked until the issue is resolved (Inc., 2023; Patel, 2022; Green, 2023; Le, 2022). For example, if a chaos experiment causes a loss of connection to external APIs, the incident management system creates a ticket assigned to the appropriate API support team, ensuring that the problem is tracked and resolved efficiently (Rivera et al., 2023; Siwach et al., 2022).
@@
This involves categorizing system components into critical and non-critical groups to focus testing efforts on the areas where failures would have the greatest impact (Dedousis et al., 2023; Anwar and Arunachalam, 2019). Critical components are essential for maintaining core operations, and their failure would cause significant disruption (Fogli et al., 2023; Kesim et al., 2020). For example, API gateways and distributed databases are critical because these failures can disrupt communication between services, cause widespread outages, or lead to data loss and downtime, directly affecting users (Siwach et al., 2022). In contrast, non-critical components have minimal immediate impact on the core functionality of the system. For instance, log aggregation systems are non-critical; their failure may affect internal monitoring but will not disrupt user-facing services or essential system functions (Dedousis et al., 2023). By prioritizing critical components, testing can focus on areas where failure would result in the most severe consequences. In contrast, the resilience issues of non-critical components can be addressed less urgently, since their failure poses a lower risk to overall system stability (Wickramasinghe, 2023).

### v2

@@ 6.1.1. Execution Environment
Virtual machines (VMs) offer a layer of abstraction over physical hardware, allowing multiple independent operating systems to share resources efficiently (Sharma et al., 2016; Smith and Nair, 2005; Basiri et al., 2016). However, VMs can experience performance issues when resources are overutilized, and they are vulnerable to hypervisor failures and migration difficulties (Azure Readiness, 2023; Fogli et al., 2023; Sousa et al., 2018). Chaos engineering addresses these challenges by testing VMs under stress conditions such as CPU or memory exhaustion to assess if VMs can recover from failures or seamlessly migrate workloads (Kostic, 2024; Chen et al., 2024; Green, 2023).
Containers, which encapsulate applications and their dependencies in lightweight environments, provide greater efficiency and portability compared to VMs (Koskinen et al., 2019; Bedoya et al., 2023). However, containers face resource contention, network instability, and orchestration failures, especially in platforms such as Docker and Kubernetes (Simonsson et al., 2021). Chaos engineering tests on containers simulate abrupt stoppages and network disruptions, helping to evaluate containers’ ability to recover within container orchestration systems while maintaining application performance  (Simonsson et al., 2021; Camacho et al., 2022; Bremmers, 2021; Kostic, 2024; Ikeuchi et al., 2020; Bedoya et al., 2023).
Serverless functions allow applications to be triggered by specific events and run in isolated containers without requiring infrastructure management (Castro et al., 2017; Taibi et al., 2020; Al-Said Ahmad et al., 2024; Aharon, 2024). While serverless architectures are scalable and cost-effective, they face challenges such as resource limitations and the reliance on third-party cloud providers (Rajan, 2020; van Eyk and Iosup, 2018; Nookala, 2023). Chaos engineering helps ensure that serverless functions can handle these challenges by simulating cold starts and execution failures, verifying that functions recover quickly and maintain performance under varying loads (Starr et al., 2022; Jernberg et al., 2020; Bedoya et al., 2023; Morichetta et al., 2023; Chatterjee, 2023).
@@
Public cloud platforms, such as AWS, Microsoft Azure, and Google Cloud, offer computing resources over the Internet that are shared between multiple organizations (Saraswat and Tripathi, 2020; Torkura et al., 2020). This model enables businesses to scale their infrastructure on demand, eliminating the need for significant investments in physical hardware. Due to its cost efficiency and scalability, the public cloud has become a popular choice for many enterprises (Dutta and Dutta, 2019; Jakkaraju, 2020; Colyer, 2019). However, relying on shared infrastructure introduces specific vulnerabilities. For example, a regional outage or failure in an availability zone can have widespread impacts, simultaneously affecting many customers (Paavolainen, 2016). In addition, many organizations depend on third-party services, such as application programming interfaces (APIs) or managed databases, and failures in these services can trigger further disruptions in their systems (Zhou et al., 2010; Krishnamurthy, 2021). To address these risks, chaos engineering can simulate large-scale failures, such as regional outages or availability zone failures, to test how well systems can failover to other regions without significant service interruptions (Fogli et al., 2023; National Australia Bank, 2020). Moreover, by injecting failures into critical third-party services, chaos engineering enables organizations to evaluate the effectiveness of their recovery strategies, ensuring that operations can continue even in the face of service disruptions (Yu et al., 2021; Hawkins, 2020; Vinisky, 2024).
@@
Private cloud environments, which provide dedicated infrastructure exclusively for a single organization, can be hosted on-premises or managed by third-party providers, offering greater control over data privacy, security, and customization (Jadeja and Modi, 2012; Gastermann et al., 2015). This model suits organizations with strict regulatory requirements or managing sensitive information (Pearson, 2013). Despite their benefits, private clouds often face scalability and geographic redundancy challenges, as they typically lack the built-in redundancy of public clouds, making them more susceptible to service interruptions due to hardware failures or resource constraints (Kumar and Goudar, 2012; Kassab et al., 2022). Without adequate failover mechanisms, private cloud environments may struggle to maintain operational continuity during hardware or network failures (Rimal et al., 2009; Tivey et al., 2019). Chaos engineering can be used in private clouds to simulate hardware failures, such as server crashes or network partitions, to evaluate the system’s ability to recover from localized disruptions (Patel, 2022; Simonsson et al., 2021). By simulating resource exhaustion, such as CPU or memory overloads, organizations can test whether their infrastructure can scale to meet unexpected surges in demand, ensuring continuous service availability under pressure (Gill, 2021; Jernberg et al., 2020).
@@
Finally, hybrid cloud environments combine the strengths of public and private clouds, allowing organizations to manage sensitive workloads on private infrastructure while leveraging the scalability of public clouds for nonsensitive or high-demand tasks (Goyal, 2014; Poltronieri et al., 2022). This approach provides flexibility but introduces complexity in maintaining smooth integration between the two environments (Lu et al., 2014). A key challenge of hybrid clouds is to ensure seamless communication between public and private components, as issues such as network latency, connectivity problems, and data synchronization delays can arise when workloads are distributed across both environments, potentially leading to service disruptions (Annapureddy, 2010; Siwach et al., 2022). Chaos engineering can address this issue by allowing organizations to simulate network interruptions or increased latency between public and private clouds (Fogli et al., 2023). These tests help to evaluate whether workloads can continue to operate despite connectivity issues (Torkura et al., 2020; Dedousis et al., 2023). Furthermore, chaos experiments that simulate delays or failures in data replication help ensure that the system maintains data consistency and operational continuity, even when part of the infrastructure is compromised (Camacho et al., 2022).
@@
This refers to the computing infrastructure hosted and managed within an organization’s physical location rather than relying on cloud providers (Willard and Hutson, 2024; Nikolaidis et al., 2021; Väisänen, 2023). These systems offer greater control over hardware, data, and network resources, essential for industries with strict regulatory requirements. However, they also face hardware failures and scaling limitations, affecting the reliability of the systems (Byström, 2022). For example, an organization managing critical on-premise data might simulate a hardware failure, such as a server or disk crash, to test how quickly backup systems respond. If synchronization delays are found, the replication process can be optimized to ensure smooth transitions and compliance with data availability standards (Fogli et al., 2023; Dedousis et al., 2023). Similarly, scaling limitations can be addressed by testing how an on-premise infrastructure responds to various demand surges, such as a sudden spike in user traffic during an e-commerce company’s high-demand event (Azure Readiness, 2023; Gianchandani et al., 2022). This helps identify bottlenecks, optimize resource allocation, and ensure that the system remains responsive under varying load conditions (Naqvi et al., 2022; Wickramasinghe, 2023; Camacho et al., 2022; Nikolaidis et al., 2023, 2021).
@@ 6.1.2. Experimentation Modes
Manual experimentation is most appropriate when engineers require high precision and control, as they take full responsibility for designing, configuring, and executing chaos tests (Kostic, 2024; Torkura et al., 2020). This approach is suited for organizations with little or no experience in chaos engineering or when highly specific, customized experiments are necessary (Durai, 2022; Zhang et al., 2019). For example, an engineer might manually trigger a database failure or introduce network latency to observe the system response in real time (Jernberg et al., 2020; Bremmers, 2021). While this mode allows for fine-tuned control, it can be time-consuming and prone to human error, making it ideal for small-scale, customized tests or critical systems where close monitoring is essential (Ikeuchi et al., 2020; Chen et al., 2024).
Semi-automated experimentation provides a middle ground between manual control and complete automation, where engineers define the experiment parameters (Jernberg et al., 2020; Pierce et al., 2021; Frank et al., 2021b). At the same time, the platform automates tasks such as triggering tests and collecting data, making this mode ideal for experiments that benefit from automated execution while still requiring moderate human involvement to reduce manual effort (Kesim et al., 2020; Nikolaidis et al., 2021; Gill, 2021). For example, engineers might set up an experiment to test resource exhaustion on specific microservices, with the platform automatically running the test at predefined intervals, collecting performance metrics, and allowing them to focus on adjusting parameters and interpreting results (Pierce et al., 2021; Dedousis et al., 2023).
Fully automated experimentation suits large-scale systems that require frequent and continuous testing (Simonsson et al., 2021; Basiri et al., 2019; Patel, 2022). After engineers configure the initial setup, the chaos engineering platform takes over, seamlessly integrating experiments into the system operations, typically through the CI/CD pipelines (Konstantinou et al., 2021; Pulcinelli et al., 2023; Manna, 2021; Colyer, 2019). For example, after a code deployment or infrastructure update, the platform can automatically simulate a service outage or introduce network disruptions, then autonomously gather performance metrics, identify system weaknesses, and generate detailed reports, requiring engineers to intervene only when necessary (Dedousis et al., 2023; Bairyev, 2023; Kesim et al., 2020; Cotroneo et al., 2022; Nikolaidis et al., 2021).
@@ 6.1.3. Chaos Automation Strategy
This involves developing programs or scripts to extend the capabilities of existing tools and gain precise control over system configurations and experiments (Myers et al., 2000; Thüm et al., 2014; Khankhoje, 2022; Cotroneo et al., 2022). It allows engineers to simulate unique failure conditions, ranging from simple component failures, such as restarting a single service, to complex multi-step sequences involving multiple systems, tailored to the specific challenges of their architecture that standard tools may not address (Gill, 2021; Inc., 2023; Torkura et al., 2020; Azure Readiness, 2023; Patel, 2022; Nikolaidis et al., 2023). For example, custom code can terminate random containers in distributed environments such as Kubernetes, simulating unexpected service failures and testing the system’s recovery capacity by automatically restarting services (Bremmers, 2021; Fawcett, 2020; Simonsson et al., 2021; Vu et al., 2022). In more complex scenarios, custom code can initiate a service outage followed by a database failure or a network partition to observe how these failures impact system recovery and resilience (Wickramasinghe, 2023; Gill, 2021; Simonsson et al., 2021). Additionally, custom code can be integrated into CI/CD pipelines to simulate high resource consumption, such as CPU throttling or memory exhaustion, after each deployment, allowing early detection of performance bottlenecks before they affect production (Azure Readiness, 2023; Konstantinou et al., 2021; Manna, 2021; Kadikar, 2023; Palani and Gupta, 2023a). This custom code for chaos experiments can be written in various programming languages, such as Python, Bash, or Go, depending on the complexity and specific requirements of the experiment (Siwach et al., 2022; Malik et al., 2023; Sondhi et al., 2021; Simonsson et al., 2021; Zhang et al., 2021b).
@@
This process automatically integrates external systems, such as monitoring, log, observability, alerting, incident management, and security platforms, into chaos engineering experiments (Torkura et al., 2020; Simonsson et al., 2021). This automated integration ensures continuous data collection without manual intervention, enabling real-time analysis during system disruptions or failures. As a result, teams can monitor performance, identify issues quickly, and respond more efficiently to failures (Palani and Gupta, 2023a; Kostic, 2024). This process is vital to integrating monitoring tools that track essential system metrics such as CPU usage, memory consumption, and network performance (Wu, 2021; Gianchandani et al., 2022; Zhang et al., 2019). Monitoring tools can automatically detect performance deviations and provide immediate feedback. For example, if a service is overloaded due to an unexpected spike in user requests, monitoring systems measure the increase in CPU usage and network load, helping the team assess the impact and determine whether the system can handle the demand or if further intervention is needed (Wickramasinghe, 2023; Fogli et al., 2023; Cotroneo et al., 2022; Gill, 2021; Siwach et al., 2022; Malik et al., 2023). Logging tools record vital events such as service failures, error messages, and recovery attempts, documenting the system’s response in real-time. For example, if network communication between two microservices fails, the logging systems capture the exact moment the failure occurred and show the error messages generated by failed retries (Cotroneo et al., 2022; Ma et al., 2023). This data helps teams analyze how the failure affected communication and whether recovery mechanisms worked effectively (Naqvi et al., 2022; Malik et al., 2023; Kostic, 2024; Palani and Gupta, 2023a). Building on this, observability tools provide a broader view by correlating logs and metrics to show how different components of the system interact during failure scenarios. These tools enable teams to identify how failures propagate across interconnected services (Simonsson et al., 2021; Manna, 2021; Green, 2023; Kostic, 2024; Krishnamurthy, 2021). For example, if latency is introduced between services in a chaos experiment, observability systems track the delays at each hop between services and how this latency affects downstream components such as database queries or API responses. This tracing allows the team to see if minor delays in one service cause issues elsewhere (Durai, 2022; Katirtzis, 2022; Kostic, 2024; Chen et al., 2024; Nikolaidis et al., 2023; Frank et al., 2023). Furthermore, alert notification systems can ensure that relevant teams are promptly notified when critical failures occur or performance thresholds are exceeded (Tivey et al., 2019; Gianchandani et al., 2022; Palani and Gupta, 2023a; Bremmers, 2021; Starr et al., 2022). For example, if a database connection pool reaches its maximum capacity during a high-load chaos test, the alert system triggers a notification to the database operations team, allowing them to address the issue before it causes a system-wide failure. This early warning helps prevent more significant disturbances (Green, 2023; Malik et al., 2023). Finally, incident management tools streamline the resolution process in conjunction with alerting by automatically generating tickets for failures detected during chaos tests. These tickets are assigned to the appropriate teams and tracked until the issue is resolved (Inc., 2023; Patel, 2022; Green, 2023; Le, 2022). For example, if a chaos experiment causes a loss of connection to external APIs, the incident management system creates a ticket assigned to the appropriate API support team, ensuring that the problem is tracked and resolved efficiently (Rivera et al., 2023; Siwach et al., 2022).
@@
This involves categorizing system components into critical and non-critical groups to focus testing efforts on the areas where failures would have the greatest impact (Dedousis et al., 2023; Anwar and Arunachalam, 2019; Chen et al., 2024). Critical components are essential for maintaining core operations, and their failure would cause significant disruption (Fogli et al., 2023; Kesim et al., 2020). For example, API gateways and distributed databases are critical because these failures can disrupt communication between services, cause widespread outages, or lead to data loss and downtime, directly affecting users (Siwach et al., 2022). In contrast, non-critical components have minimal immediate impact on the core functionality of the system. For instance, log aggregation systems are non-critical; their failure may affect internal monitoring but will not disrupt user-facing services or essential system functions (Dedousis et al., 2023). By prioritizing critical components, testing can focus on areas where failure would result in the most severe consequences. In contrast, the resilience issues of non-critical components can be addressed less urgently, since their failure poses a lower risk to overall system stability (Wickramasinghe, 2023; Inc., 2023).